OpenStudy (underscorerocky):

How can 1/5x – 2 = y and 1/3x + 8 = y be set up as a system of equations? (5 points) a. 5y – 5x = –10 3y – 3x = 24 b. 5y – 5x = –10 3y + 3x = 24 c. 5y + x = –10 3y + x = 24 d. 5y – x = –10 3y – x = 24

OpenStudy (underscorerocky):

I will medal anyone who solves this correctly

OpenStudy (thadds2003):

Ok, hold on.

OpenStudy (thadds2003):

How many answers are there?

OpenStudy (thadds2003):

Sorry, this is a test and i'm not allowed to help on tests..... hears a medal anyway.

OpenStudy (underscorerocky):

@Thadds2003 I already answered it I just want to see how it is solved

OpenStudy (underscorerocky):

@perl @Michele_Laino can you show me?

jhonyy9 (jhonyy9):

how is right ? 1 ---x-2 = y 5 1 --x+8 = y 3 in this way is correct wrote ?

jhonyy9 (jhonyy9):

than right so will get x-5*2 =5y x+8*3=3y rearranging them like a system will get x-5y=10 x-3y=-24 hope so much that this will help you

OpenStudy (underscorerocky):

it is written as 1/5x-2=y and 1/3x+8=y, i dont know if that changes your answer or not.

jhonyy9 (jhonyy9):

so this mean that 5x-2 and 3x+8 are denominators ?

OpenStudy (underscorerocky):

yes, sorry I should write it as (1/5)x-2 and (1/3)x+8

OpenStudy (underscorerocky):

it doesnt mean that 5x-2 and 3x+8 are denominators my bad.

jhonyy9 (jhonyy9):

so this mean that ,,not" is right how i have thought it and how i have made my calculi there above a. 5y – 5x = –10 3y – 3x = 24 b. 5y – 5x = –10 3y + 3x = 24 c. 5y + x = –10 3y + x = 24 d. 5y – x = –10 3y – x = 24 x-5y=10 than rewrite this -5y+x=10 and multiplie both sides by -1 what will get x-3y=-24 and same this too like the first - try it please and check the choice d.

OpenStudy (underscorerocky):

@jhonyy9 thank you so much. sorry for making it harder than it had to be lol.

jhonyy9 (jhonyy9):

was my pleasure anytime than i m here good luck bye

jhonyy9 (jhonyy9):

so in the next time use please parentheses to be more easy understandably - ok. ?